Please help me . This is due tomorrow so I would really apriciate to get the answer ASAP.
A watch, set accurately at noon,at 5p.m indicates 10min. to 5.How many minutes past 5 is it actualyy when the watch indicates 5p.m.. Express your answer as a mixed number.

Answer by josmiceli(19441)

The watch loses 10 minutes during 5 hours, so it's
losing 10 minutes every 300 minutes. It's time indicated is
T(watch) = (290/300)*T(accurate)
300 = (290/300)*Ta
Ta = 9000/290
Ta = 310 10/29
So 10 10/29 is the minutes past 5pm the watch reads when it's
actually 5pm